OpenStudy (anonymous):

Find the sum of the following infinite geometric series, if it exists. 2 + 1.5 + 1.125 + 0.8437 +…

OpenStudy (anonymous):

Find the sum of the infinite series 3 + 1.2 + 0.48 + 0.192 + …if it exists.

OpenStudy (anonymous):

use the a/1-r equation again.. your a=3 and your r = 2.5. Because to get from 3 to 1.2 you divide 3 by 2.5 or (5/2). I believe you should get 2.

OpenStudy (anonymous):

r =.75 so sum of serease= a/(1-r) 2/(1-.75)=8 sum of series =8

OpenStudy (anonymous):

oh geez ignore my comment. wrong one

OpenStudy (anonymous):

sam to what question have answered

OpenStudy (anonymous):

sam's is to the 1st question. mine is to the second one.

OpenStudy (anonymous):

okay...same one r=.4 a=3 so sum of series=3/(1-.4)=5